OpenStudy (anonymous):

Find the equation of the circle that satisfies the conditions: Center at the origin, pass through P(4,-7)

OpenStudy (anonymous):

I got \[x^2+y^2=65\]

OpenStudy (lgbasallote):

the radius would be distance formula from point (4,-7) to (0,0)

OpenStudy (anonymous):

Yes! You got it!

OpenStudy (amistre64):

r^2 = x^2 + y^2 :)

OpenStudy (amistre64):

no, h and k are the origin 0,0

OpenStudy (lgbasallote):

@ParthKohli center at origin

OpenStudy (amistre64):

4,-7 is the point that the radius hits

OpenStudy (anonymous):

Thanks everyone!!!

OpenStudy (lgbasallote):

it's good that you're improving @purplec16 ^_^

OpenStudy (anonymous):

huh?

OpenStudy (amistre64):

uhoh, it appears that in order to learn math better she might of had to forgotten some english skills :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

lol...